OpenStudy (anonymous):

f(x) = (x^2 +2)(x-3) is the f'(x) = 3x^2 -6x +2?

jimthompson5910 (jim_thompson5910):

you are correct again

OpenStudy (anonymous):

(x^2 + 2)(x - 3) = x^3 - 3x^2 + 2x - 6 = f(x) d(f(x))/dx = 3x^2 - 6x + 2